OpenStudy (twalt13):

Factor the algebraic expression below in terms of a single trigonometric function. sin^2 x +sin x-2

OpenStudy (anonymous):

\[( \sin(x)-1 )(\sin(x)+2)\]

OpenStudy (anonymous):

substitute sin(x) with (y)

OpenStudy (mathstudent55):

Let u = sin x Then sin^2 x + sin x - 2 = u^2 + u - 2 u^2 + u - 2 is easy to factor: (u - 1)(u + 2) Now substitute back sin x for u: (sin x - 1)(sin x + 2)

OpenStudy (anonymous):

you will get : \[ y^2 + y -2\]

OpenStudy (anonymous):

now find two # Multiplied = -2 and the SUM of them = +1

OpenStudy (anonymous):